# First Hitting Times on a Quantum Computer: Topological Effects and Dark States

### 24 Feb 2024 v1

@@ VI.3 Tracking protocol (broadening)
Close to the transition $\langle n\rangle=1$ for any large though finite $N$. If we choose a control parameter like $\gamma\tau$ to be far from the transition, then $\langle n\rangle=3$. Fig. 8 compares the approximations for the on-site and tracking protocol in Eq. 13 and Eq. 16 with the simulation and experimental data.
@@
VI.4 Slow decay of the null measurement probability
@@
It is zero in the limit $N\rightarrow\infty$ for the return problem [5, 8]. Near special points where the three phase factors merge, the null measurement probabilities decay only very slowly to zero for increasing $N$, leading to the transition of $\langle n\rangle$ to 3 (tracking protocol) and 2 (on-site protocol) only for large $N>1000$ for $\gamma\tau=2\pi/(\sin(0.5+\pi/6)+\cos(0.5))$ [Fig. 7 (c)]. If $\gamma\tau\approx 3.63$, $\langle n\rangle$ shows a staircase behavior. The system is only partially thermalized until, for a larger number of measurements, the high-temperature limit is reached.
@@ VII Conclusion
Hitting times and detection probabilities are closely related to the number of distinct eigenvalues (phase factors) of the unitary time-evolution operator $U=\exp(-iH\tau)$ of the model Hamiltonian $H$ during the measurement-free evolution time $\tau$ shown in Fig. 2. To investigate them, we swept through the parameters of the phase-factor matching diagram and performed two different stroboscopic measurement protocols. In the first protocol, readout was performed only at the target site. The second protocol, known as the tracking protocol, involves measurements at each site along the ring. The main experimental result is that there are clear differences between both protocols, a feature not known for classical random walks. The results show that the postulates of measurement theory could be confirmed on the IBM quantum computer, a non-trivial fact since $N$ repeated noisy measurements are utilized.
@@
We showed that finite resolution effects, which are clearly part of any experimental study, are a key feature of the hitting time statistics. This leads to a broadening of the mean return time and the detection probability because the detection is imperfect near special points, where two or three phase factors merge, due to finite $N\tau$. The finite resolution and the broadening effect are accompanied by a slow relaxation of the null measurement probability. The results show not only a broadening of the topological transitions but also metastable topological [see Fig. 3 (e) and see Fig. 4 (e)] and chirality (see Fig. 6) effects, which disappear for the theory in the asymptotic limit.
@@
We derived an analytical expression for the broadening of the $\langle n\rangle=2$ to $\langle n\rangle=1$ and back transition (on-site) and the $\langle n\rangle=3$ to $\langle n\rangle=1$ and back transition (tracking) for $\alpha=0$ as a function of $N$. For both protocols, the resonances narrow as $N$ increases. Furthermore, the additional resonances near the special points (but not exactly on them) disappear for growing $N$. The mean return time $\langle n\rangle$ shows a crossover between different topological phases, which is revealed, e.g., as the staircase behavior for the on-site protocol [Fig. 7 (c)]. A similar effect was found for a perturbed ring [14]. The tracking protocol of the monitored quantum walk has more of the character of a classical random walk. Because of the finite resolution, quantum effects are present for a small number of midcircuit measurements close to the revival time. This can be seen as partial thermalization [37] during a quantum-to-classical crossover. Only for a large number of measurements the classical limit and maximum $\langle n\rangle$ is reached.
@@
There are several future directions in which our work can be extended. Here, we focus on small systems. More generally, one could scale up the process by increasing the number of measurements and/or increasing the size of the system. One could then ask whether the quantum computer exhibits a transition to a more classical behavior, for example, the elimination of the topological effect found for the mean return time. Furthermore, one could explore the possibility of measuring the readout rate-controlled oscillatory behavior of the energy expectation values corresponding to a certain trajectory of the monitored dynamics. In a future publication, we will show that noise in the system enables transitions to theoretically forbidden states, an effect that becomes particularly important as we increase $N$.

### 10 Apr 2024 v2

@@ VI.1.2 Tracking protocol (broadening)
Close to the transition $\langle n\rangle=1$ for any large enough $N$. If we choose a control parameter like $\gamma\tau$ to be far from the transition, then $\langle n\rangle=3$. Fig. 8 compares the approximations for the on-site and tracking protocol in Eq. 13 and Eq. 16 with the simulation and experimental data.
@@
VI.2 Slow decay of the null measurement probability
@@
It is zero in the limit $N\rightarrow\infty$ for the return problem [5, 6, 9]. Near special points where the three phase factors merge, the null measurement probabilities decay only very slowly to zero for increasing $N$, leading to the transition of $\langle n\rangle$ to 3 (tracking protocol) and 2 (on-site protocol) only for large $N>1000$ for $\gamma\tau=2\pi/(\sin(0.5+\pi/6)+\cos(0.5))$ [Fig. 7 (c)]. If $\gamma\tau\approx 3.63$, $\langle n\rangle$ shows a staircase behavior. The system is only partially thermalized until, for a larger number of measurements, the high-temperature limit is reached.
@@ VII Summary and Conclusion
Hitting times and detection probabilities are closely related to the number of distinct eigenvalues (phase factors) of the unitary time-evolution operator $U=\exp(-iH\tau)$ of the model Hamiltonian $H$ during the measurement-free evolution time $\tau$ shown in Fig. 2. To investigate them, we swept through the parameters of the phase-factor matching diagram and performed two different stroboscopic measurement protocols. In the first protocol, readout was performed only at the target site. The second protocol, known as the tracking protocol, involves measurements at each site along the ring. The main experimental result is that there are clear differences between both protocols, a feature not known for classical random walks. The onsite protocol shows quantum behavior. The mean return time is quantized and linked to a winding number. Moreover, the winding number of the return problem is equal to the dimensionality of the available Hilbert space. This was originally discovered in Ref. [5, 6] for periodic measurements and later found also for measurements at random times [17]. Since the dimensionality of the available Hilbert space changes when degenerate states occur, the winding number as well as the mean return time can jump between discrete values.
The tracking protocol exhibits quantum features when all the phase factors match, but otherwise shows classical behaviors. The results show that the postulates of measurement theory as well as the presence of a phase utilising the onsite protocol could be confirmed on the IBM quantum computer, a non-trivial fact since $N\sim 20$ repeated noisy measurements are utilized.
@@
We showed that finite resolution effects, which are clearly part of any experimental study, are a key feature of the hitting time statistics. This leads to a broadening of the mean return time and the detection probability because the detection is imperfect near special points, where two or three phase factors merge, due to finite $N\tau$. The finite resolution and the broadening effect are accompanied by a slow relaxation of the null measurement probability. The results show not only a broadening of the topological transitions but also metastable topological [see Fig. 3 (e) and see Fig. 4 (e)] and chirality effects (see Fig. 6), which disappear for the theory in the asymptotic limit.
@@
We derived an analytical expression for the broadening of the $\langle n\rangle=2$ to $\langle n\rangle=1$ and back transition (on-site) and the $\langle n\rangle=3$ to $\langle n\rangle=1$ and back transition (tracking) for $\alpha=0$ as a function of $N$. For both protocols, the resonances narrow as $N$ increases. Furthermore, the additional resonances near the special points (but not exactly on them) disappear for growing $N$. The mean return time $\langle n\rangle$ shows a crossover between different topological phases, which is revealed, e.g., as the staircase behavior for the on-site protocol [Fig. 7 (c)]. A similar effect was found theoretical for a perturbed ring [15]. As mentioned, the tracking protocol of the monitored quantum walk has more of the character of a classical random walk. Because of the finite resolution, quantum effects are present close to the revival time. This can be seen as partial thermalization [41] during a quantum-to-classical crossover. Only for a large number of measurements the classical limit and maximum $\langle n\rangle$ is reached, except for special $\gamma\tau$ and $\alpha$.
@@
There are several future directions in which our work can be extended. Here, we focus on small systems. More generally, one could scale up the process by increasing the number of measurements and/or increasing the size of the system. One could then ask whether the quantum computer exhibits a transition to a more classical behavior, for example, the elimination of the topological effect found for the mean return time or how a partial information about the trajectories by measuring multiple but not all sites change the topological effect. Furthermore, one could explore the possibility of measuring the readout rate-controlled oscillatory behavior of the energy expectation values corresponding to a certain trajectory of the monitored dynamics. In a future publication, we will show that noise in the system enables transitions to theoretically forbidden states, an effect that becomes particularly important as we increase $N$.